OpenStudy (anonymous):

solve 2cos^2 x - cosx - 1 = 0 where 0

OpenStudy (anonymous):

2cos^2 x - cosx - 1 = 0 2cos^2 x - 2cosx +cosx- 1 = 0 2cosx (cosx-1) + 1(cosx -1) =0 (cosx-1)(2cosx+1) =0 cosx =1,-1/2 hope u can do further..

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw